OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

3x - y = ? Parallel lines have the same slope so just plug in the coordinate. 3(4) - (-2) = 14 3x - y = -3 This is the same question as before.